# RoboMemory: A Brain-inspired Multi-memory Agentic Framework for Lifelong Learning in Physical Embodied Systems

### v3

@@ 3.2.2 Lifelong Learning System
In real-world scenarios, agents process sequential tasks over their lifetime, requiring continuous improvement through prior experiences. Drawing on cognitive psychology’s classification of human long-term memory, we divide our system into episodic and semantic memory: the former records agent-environment interaction histories, while the latter extracts experiential insights to support long-term task reasoning. This update process mirrors human daily experience consolidation during sleep maboudi2024retuning .
@@ 3.3 Closed-Loop Planning Module for Dynamic Environment
To enable closed-loop control in embodied environments, the Closed-Loop Planning Module adopts the Planner-Critic mechanism lei2025clea , which consists of the planner and the critic module. For each planning step, the planner generates a long-term plan consisting of multiple steps. However, due to the dynamics of embodied environments, the action sequence in the long-term plan may become outdated during the execution of the plan. Thus, before executing each step, we use the Critic model to evaluate whether the proposed action in this step remains appropriate under the latest environment. If not, the planner will replan based on the latest information. The demonstration of this process is shown in Fig. 2.
@@ 3.4 Low-level Executer
We use $\pi_{0}$ black2024pi_0 as the action executor in the real world: actions planned by RoboMemory are converted into arm and chassis movements via the Low-level Executer, with LoRA fine-tuning applied to optimize performance in real-world robotic tasks. More details are provided in Appendix B.
@@ 4 Experiments
4.1 Experimental Setup
4.1.1 Benchmarks
To evaluate the task planning ability of RoboMemory, we select a subset of the EB-ALFRED benchmark from EmbodiedBenchyang2025embodiedbench . We selected the Base and Long subsets because they aim to test the agent’s planning ability. The Base and Long subset comprises 100 tasks for complex long-term embodied tasks. The EB-ALFRED environment provides a visually-grounded operational setting that closely mimics real-world conditions (see appendix B for environment details), enabling direct comparison with established baselines.
Additionally, we have also evaluated RoboMemory’s capabilities on the EB-Habitat benchmark yang2025embodiedbench . For detailed results, please refer to Appendix A.
Moreover, we set up a real-world environment to test the lifelong learning ability of RoboMemory in the real world.
4.1.2 Settings & Baselines
To facilitate comparisons, we consider two types of baselines. First, we choose the advanced closed-source and open-source VLMs as a single agent. We compare their performance with RoboMemory. For closed source VLMs, we choose GPT-4o and GPT-4o-mini GPT-4o-mini ; hurst2024gpt , Claude3.5-Sonnet Claude-3.5-Sonnet , Gemini-1.5-Pro and Gemini-2.0-flash team2024gemini ; Gemini2.0 . For open source VLMs, we choose LLama-3.2-90B-Vision-Ins llama3.2 , InternVL-2.5-78B/28B chen2024expanding , InternVL-3-72B zhu2025internvl3 , and bai2025qwen2 . Secondly, we choose three agent frameworks: (1) Reflexion shinn2024reflexion , which introduces a simple long-term memory and a self-reflection module. Reflexion uses the self-reflection module to summarize experiences as long-term memory, thereby enhancing the model’s capabilities. (2) Voyager wang2023voyager , which utilizes a skill library as its procedural memory, is a widely used baseline for embodied agent planning. (3) Cradle tan2024cradle , which proposes a general agent framework with episodic and procedural memory and gains good performances at various multi-model agent tasks.
In our experiments, each agent framework is tested using Qwen2.5-VL-72b-Ins qwen2.5 . The Qwen2.5-VL-72b-Ins represents a high-performing open-source alternative. Notably, the Qwen2.5-VL-72b-Ins demonstrates performance comparable to advanced closed-source VLMs in several benchmark tasks white2024livebench . We use the Qwen3-Embedding model zhang2025qwen3 to create embedding vectors for RAGs in RoboMemory. For the Low-level Executor, since EB-ALFRED provides high-level action APIs, we use the low-level executor provided by EmbodiedBench instead of the VLA-based method.
4.1.3 Evaluation metrics
@@
Figure 3: The reason why RoboMemory failed to complete the task.
@@ 4.2 Main results
As shown in Table 1, our model achieves significant improvements over both single VLM agents and Agent frameworks on the EB-ALFRED. Compared to the SOTA Single VLM-Agent model, Claude3.5-Sonnet, RoboMemory with Qwen2.5-VL-72B-Ins backbone improves the average SR by 5% and GC by 15%. This demonstrates RoboMemory’s superiority over single VLM-Agents, proving that an Agent framework with open-source models can outperform closed-source SOTA models. Furthermore, when tested against other VLM-Agent frameworks, RoboMemory also shows substantial gains. This is because, unlike other frameworks, RoboMemory’s brain-like memory system provides embodied models with more accurate and persistent contextual information. Additionally, the Planner-Critic mechanism provides a closed-loop planning ability, which helps the RoboMemory gain better performance in long-term tasks. Because the RoboMemory can detect and try to overcome possible failures. And it is more robust when encountering unexpected situations.
@@ 4.3 Ablation Studies
4.3.1 Long-term Memory
Adding long-term memory significantly improved RoboMemory’s success rate. The experiment shows that it enables continuous learning while it attempts to complete tasks. The semantic memory learns low-level skills’ properties, such as in what circumstances an action may fail. The temporal memory records all task attempts (successful/failed), providing valuable experience at the task level, giving insight about how to complete a task successfully. This helps the RoboMemory predict action outcomes and avoid ineffective attempts. This ability indicates that the RoboMemory has lifelong learning capability.
4.3.2 Spatial Memory
Spatial memory is crucial for embodied agents, especially given that current pretrained VLMs have limited spatial understanding ability. Our novel dynamic KG update algorithm enables KG-based spatial memory in dynamic environments. This spatial reasoning helps RoboMemory handle partially observable embodied settings.
4.3.3 Critic Module
Table 2 shows performance without the critic module (55% vs 67% with full system). This drop highlights how the critic’s closed-loop planning adapts to dynamic environments. It helps RoboMemory recover from failures faster and handle unexpected situations better.
4.4 Error Analysis
We analyze RoboMemory trajectories for failed tasks. We identify error types based on the above definitions. A single task may contain multiple errors. We calculate the occurrence probability of each error type to show RoboMemory’s strengths and weaknesses. The results are shown in Fig. 3.
The detailed examples and discussions are provided in Appendix D.
@@ 4.5 Real-world Robot Deployment
4.5.1 Environmental Setup
To evaluate RoboMemory’s lifelong learning capability in the real world, we designed a kitchen environment inspired by EB-ALFRED and EB-Habitat. The scene contains 5 navigable points, 8 interactive objects, and over 10 non-interactive (but potentially distracting) items. The environment is shown in Fig. 5. In the real world, we use continuous video recordings during action execution (instead of snapshots after the action execution) as RoboMemory’s input. This provides a more temporally coherent perception. We created three task categories (5 tasks each, totaling 15 tasks) with difficulty matching EB-ALFRED’s Base subset (average oracle trajectory length from 5 to 10). Additional hardware experiment details are in Appendix B.
To test the lifelong learning ability of RoboMemory, we run each task twice without clearing long-term memory between attempts. Success rates for first and second attempts are shown in Fig. 5.
4.5.2 Main results
The second attempt showed significantly higher success rates. This proves RoboMemory’s long-term memory effectively guides subsequent tasks in real embodied environments. Key observations include:
  1. 1.
    Closed-loop error recovery: RoboMemory retries failed actions when possible, even if the low-level executor (VLA model) fails.
  2. 2.
    Spatial reasoning: RoboMemory remembers object locations and spatial relationships using its memory.
  3. 3.
    Lifelong learning: RoboMemory analyzes failure causes reasonably. These analyses guide future decisions.
Detailed examples demonstrating these capabilities and further discussions are provided in Appendix D.
@@ 5 Conclusion and Future Work
In conclusion, RoboMemory, a brain-inspired multi-memory framework, enables lifelong learning and long-term planning in real-world embodied systems by addressing key challenges: continuous learning, memory latency, task correlation capture, and planning infinite loops. Experimental results on EmbodiedBench show it outperforms SOTA closed-source VLMs and agent frameworks, with ablation studies validating critical components like the Critic module and spatial/long-term memory. Real-world deployment confirms its lifelong learning capability via improved success rates in repeated tasks. While limited by reasoning errors and executor reliance, it paves the way for generalizable memory-augmented agents, with future work focusing on refining reasoning and enhancing execution robustness.
A key unsolved problem in current hierarchical agent research for embodied tasks, including ours, is about the interaction between high-level agents and low-level executors (e.g., VLAs). Most existing frameworks use language instructions solely as action instructions from high-level agents. However, some action details are hard to describe with language. Other modalities (e.g., vision) can better represent these details (e.g., grasp points). While our work focuses on long-term planning and lifelong learning for agents, future work can consider enhancing generalization within existing frameworks by improving the interaction method between the VLA and the Agent.
Appendix A Additional Experiments
@@
Figure 6: Comparison of Success Rates (SR) and Goal Condition Success Rates (GC) across difficulty levels between RoboBrain and baseline methods on EB-Habitat.
A.1 Evaluate on EB-Habitat
@@
Similar to EB-ALFRED, we also deploy RoboMemory to EB-Habitat. Both EB-Habitat and EB-HALFRED are subsets of EmbodiedBench. We evaluate our model on the Base and Long subsets. Each subset contains 50 different trails. The results are shown in Fig 6.

### v4

@@ 2.2.2 Interactive Environmental Learning System
Although episodic memory can record “what happened,” it cannot directly answer “why it succeeded/failed” or “what should be done next time.” Drawing on cognitive psychology’s classification of human long-term memory, we divide our system into episodic and semantic memory: the former records agent-environment interaction histories, while the latter extracts experiential insights to support long-term task reasoning. This update process mirrors human daily experience consolidation during sleep (maboudi2024retuning, ).
@@ 2.3 Closed-Loop Planning Module for Dynamic Environment
To enable closed-loop control in embodied environments, the Closed-Loop Planning Module adopts the Planner-Critic mechanism (lei2025clea, ), which consists of the planner and the critic module. For each planning step, the planner generates a long-term plan consisting of multiple steps. However, due to the dynamics of embodied environments, the action sequence in the long-term plan may become outdated during the execution of the plan. Thus, before executing each step, we use the Critic model to evaluate whether the proposed action in this step remains appropriate under the latest environment. If not, the planner will re-plan based on the latest information. The demonstration of this process is shown in Figure 2.
@@ 2.4 Low-level Executor
We employ a LoRA-finetuned VLA model, $\pi_{0}$ (hu2022lora, ; black2024pi_0, ), to generate manipulation actions, and a SLAM-based navigation model for locomotion. The low-level executor then translates high-level actions planned by RoboMemory into concrete arm and chassis movements in the real world.
@@ 3 Experiments
3.1 Benchmarks
To evaluate the task planning ability of RoboMemory, we select a subset of the EB-ALFRED benchmark from EmbodiedBench(yang2025embodiedbench, ). We selected the Base and Long subsets because they aim to test the agent’s planning ability. The Base and Long subset comprises 100 tasks for complex embodied tasks. The EB-ALFRED environment provides a visually-grounded operational setting that closely mimics real-world conditions (see Appendix D for environment details), enabling direct comparison with established baselines.
Additionally, we have also evaluated RoboMemory’s capabilities on the EB-Habitat benchmark (yang2025embodiedbench, ). For detailed results, please refer to Appendix B.
Moreover, we set up an environment to test the interactive environmental learning ability of RoboMemory in the real world.
3.2 Settings & Baselines
To facilitate comparisons, we consider two types of baselines. First, we choose the advanced closed-source and open-source VLMs as a single agent. We compare their performance with RoboMemory. For closed source VLMs, we choose GPT-4o and GPT-4o-mini (GPT-4o-mini, ; hurst2024gpt, ), Claude3.5-Sonnet (Claude-3.5-Sonnet, ), Gemini-1.5-Pro and Gemini-2.0-flash (team2024gemini, ; Gemini2.0, ). For open source VLMs, we choose Llama-3.2-90B-Vision-Ins (llama3.2, ), InternVL-2.5-78B/28B (chen2024expanding, ), InternVL-3-72B (zhu2025internvl3, ), and Qwen2.5-VL-72B-Ins (bai2025qwen2, ). Secondly, we choose three agent frameworks: (1) Reflexion (shinn2024reflexion, ), which introduces a simple long-term memory and a self-reflection module. Reflexion uses the self-reflection module to summarize experiences as long-term memory, thereby enhancing the model’s capabilities. (2) Voyager (wang2023voyager, ), which utilizes a skill library as its procedural memory, is a widely used baseline for embodied agent planning. (3) Cradle (tan2024cradle, ), which proposes a general agent framework with episodic and procedural memory and gains good performances at various multi-model agent tasks.
In our experiments, each agent framework is tested using Qwen2.5-VL-72b-Ins (qwen2.5, ). The Qwen2.5-VL-72b-Ins represents a high-performing open-source alternative. Notably, the Qwen2.5-VL-72b-Ins demonstrates performance comparable to advanced closed-source VLMs in several benchmark tasks (white2024livebench, ). We use the Qwen3-Embedding model (zhang2025qwen3, ) to create embedding vectors for RAGs in RoboMemory. For the Low-level Executor, since EB-ALFRED provides high-level action APIs, we use the low-level executor provided by EmbodiedBench instead of the VLA-based method.
@@ 3.2 Settings & Baselines
Figure 3: Efficiency improvement of Comprehensive Embodied Memory System
@@ 3.3 Main results
As shown in Table 1, our model achieves significant improvements over both single VLM agents and Agent frameworks on the EB-ALFRED. Compared to the SOTA Single VLM-Agent model, Gemini1.5-Pro, RoboMemory with Qwen2.5-VL-72B-Ins backbone improves the average SR by 3% and GC by 8.7%. This demonstrates RoboMemory’s superiority over single VLM-Agents, proving that an Agent framework with open-source models can outperform closed-source SOTA models. Furthermore, when tested against other VLM-Agent frameworks, RoboMemory also shows substantial gains. This is because, unlike other frameworks, RoboMemory’s brain-like memory system provides embodied models with more accurate and persistent contextual information. Additionally, the Planner-Critic mechanism provides a closed-loop planning ability, which helps the RoboMemory gain better performance in long-term tasks. Because the RoboMemory can detect and try to overcome possible failures. And it is more robust when encountering unexpected situations.
3.4 Efficiency Analysis
To evaluate the efficiency of the Comprehensive Embodied Memory module, we tasked RoboMemory with executing 10 long-horizon tasks, each comprising approximately 50 steps. We exclusively measured the wall-clock time consumed by memory update and retrieval operations. We analyzed the scaling behavior of memory update latency across three distinct configurations: (1) fully parallel update and retrieval across all memory modules; (2) sequential update of each memory module without parallelism; and (3) update of only the most fundamental memory component: the Temporal Memory. Results are presented in Figure 3. As shown, our parallel update strategy enables updating a multi-module memory system with latency comparable to that of updating a single base memory module. This demonstrates the critical efficiency gains afforded by parallelization across the memory architecture.
@@ 3.5 Ablation Studies
Long-term Memory Adding long-term memory significantly improved RoboMemory’s success rate. The experiment shows that it enables interactive environmental learning while attempting to complete tasks. The semantic memory learns low-level skills’ properties, such as in what circumstances an action may fail. The temporal memory records all task attempts (successful/failed), providing valuable experience at the task level and giving insight into how to complete a task successfully. This helps the RoboMemory predict action outcomes and avoid ineffective attempts. This ability indicates that the RoboMemory has an interactive environmental learning capability.
Spatial Memory Spatial memory is crucial for embodied agents, especially given that current pretrained VLMs have limited spatial understanding ability. Our novel dynamic KG update algorithm enables KG-based spatial memory in dynamic environments. This spatial reasoning helps RoboMemory handle partially observable embodied settings.
Critic Module Table 2 shows performance without the critic module (55% vs 67% with full system). This drop highlights how the critic’s closed-loop planning adapts to dynamic environments. It helps RoboMemory recover from failures faster and handle unexpected situations better.
@@ 3.6 Real-world Robot Deployment
To evaluate RoboMemory’s interactive environmental learning capability in the real world, we designed a kitchen environment inspired by EB-ALFRED and EB-Habitat. The scene contains 5 navigable points, 8 interactive objects, and over 10 non-interactive (but potentially distracting) items. The environment is shown in Figure 5. In the real world, we use interactive environmental video recordings captured during action execution (rather than static snapshots taken after action completion) as RoboMemory’s input. This provides a more temporally coherent perception. We created three task categories (5 tasks each). Tasks are matched to EB-ALFRED’s Base subset (avg. oracle: 10–20 steps), though actual executions often exceed 20 steps due to search and error recovery. Due to search and error recovery, the robot often exceeds 20 steps per task. Additional hardware experiment details are in Appendix D.
To test the interactive environmental learning ability of RoboMemory, we run each task twice without clearing long-term memory between attempts. Success rates for first and second attempts are shown in Figure 5.
The second attempt showed significantly higher success rates. This proves RoboMemory’s long-term memory effectively guides subsequent tasks in real embodied environments. Key observations include: (1) Closed-loop error recovery: RoboMemory retries failed actions when possible, even if the low-level executor (VLA model) fails. (2) Spatial reasoning: RoboMemory remembers object locations and spatial relationships using its memory. (3) Interactive environmental learning: RoboMemory analyzes failure causes reasonably. These analyses guide future decisions. Detailed examples demonstrating these capabilities and further discussions are provided in Appendix E.
@@ 4 Conclusion and Future Work
In summary, RoboMemory, a brain-inspired multi-memory framework, facilitates long-horizon planning and interactive environmental learning in real-world embodied systems by addressing key challenges such as memory latency, task correlation capture, and planning loops. Experiments on EmbodiedBench demonstrate that RoboMemory outperforms state-of-the-art closed-source VLMs and agent frameworks, with ablation studies confirming the critical roles of the Critic module and spatial/long-term memory. Real-world deployment further validates its interactive learning capability through improved success rates in repeated tasks. Despite limitations arising from reasoning errors and executor dependence, RoboMemory provides a foundation for generalizable, memory-augmented agents, with future work aimed at refining reasoning and enhancing execution robustness.
A notable open challenge in hierarchical embodied agents, including RoboMemory, lies in the interface between high-level planners and low-level executors. Existing frameworks typically rely on language instructions to convey actions, yet some execution details (e.g., precise grasp points) are difficult to describe textually and are better captured through other modalities, such as vision. While our current work emphasizes long-term planning and interactive learning, future research may improve generalization by developing richer multimodal interactions between the agent and executor.
Appendix A Related Work
A.1 VLM/LLM-based Agentic Frameworks in Embodied Tasks
The rapid advancement of VLMs/LLMs has led to diverse agent frameworks in embodied environments [42, 33, 24]. Embodied tasks involve partial observability and long-horizon planning, requiring memory systems to retain context. Some use time-ordered context buffers for short-term memory (due to VLMs/LLMs’ limited long-context processing) [42, 29]; others adopt experience buffers as long-term semantic memory [14, 32]. For long-duration tasks, skill libraries serve as procedural memory, with agents accumulating skills via interaction [38, 35]. However, in real-world settings, the low-level executor may fail to complete the task, making it challenging to construct a reusable, code-based skill library. So, explicit procedural memory still needs to be improved in real-world settings. Moreover, Recent efforts integrate diverse memories [44, 35, 1] but focus on virtual/GUI environments, leaving real-world multi-modal memory support for long-term planning under-explored.
A.2 Vision Language Action Model
Current work on VLA models uses imitation learning to output low-level controls from language and visuals [5, 47, 4, 22] but is limited to tabletop tasks and single actions, restricting long-horizon planning. VLAs lack long-term execution abilities, while high-level agents excel at planning. Recent works combine high-level frameworks with VLA executors, some augmented with simple memory [31, 34, 43, 41] for longer tasks. However, real-world robots need more sophisticated memory to handle continuous multi-task operations over extended periods.
A.3 Memory Frameworks
Many previous works improve long-term planning via memory systems: Voyager [38] uses a skill library in Minecraft but lacks diverse memory types; CoELA [44] includes procedural, semantic, and episodic memory with a task-specific 2D map; MSI-Agent [14] utilizes insight as long-term memory for in-task learning. Hippo Retrieval Augmented Generation (RAG) [18] mimics the hippocampus and introduces KGs as long-term memory indices [6, 9], enhancing retrieval. However, the previous approach is mainly focused on constructing a KG with a static long context, such as a book, but it is hard to update the graph. We need to update the information in KG for the embodied task. Our approach builds a more general LLM-based memory system using a dynamic KG like Hippo RAG, which is designed for embodied tasks. Furthermore, we summarize the differences among different memory systems in previous work. The comparison is shown in the Table 3.
Table 3: Comparison of Memory-Related Methods in Embodied Agent
Method Multimodal Episodic Semantic Spatial Temporal Procedural Memory Implementation Real Robot
NeSyC [12] Symbolic logic rules
Reflexion [32] Buffer
Voyager [38] RAG
MSI-Agent [14] Database, RAG
CoELA [44] Top-down semantic map
Cradle [35] RAG
Agent-S [1] RAG
Expel [46] Buffer
AutoManual [8] Buffer
HiRobot [31] /
Being-0 [43] Buffer
RoboOS [34] Scene graph, database
RoboMemory (Ours) RAG,KG
Appendix B Additional Experiments
B.1 Error Analysis
Figure 6: The reason why RoboMemory failed to complete the task.
We analyze RoboMemory trajectories for failed tasks. We identify error types based on the above definitions. A single task may contain multiple errors. We calculate the occurrence probability of each error type to show RoboMemory’s strengths and weaknesses. The results are shown in Figure 6.
The detailed examples and discussions are provided in Appendix E.
B.2 Evaluate on EB-Habitat
@@
Figure 7: Comparison of Success Rates (SR) and Goal Condition Success Rates (GC) across difficulty levels between RoboMemory and baseline methods on EB-Habitat.
@@
Similar to EB-ALFRED, we also deploy RoboMemory to EB-Habitat. Both EB-Habitat and EB-HALFRED are subsets of EmbodiedBench. We evaluate our model on the Base and Long subsets. Each subset contains 50 different trails. The results are shown in Figure 7.
@@
B.3 Additional Efficiency analysis
We analyze the evolution of the spatial KG during long trajectories in EB-ALFRED, focusing on the first 20 iterations (with 95% confidence intervals). As shown in Figure 8, the total number of spatial relationships in the KG (red line) increases gradually over iterations as RoboMemory is exploring the environment. In contrast, the number of relationships retrieved for update at each iteration (blue line) remains relatively stable, typically ranging around 10 edges per iteration. This stability is achieved because our method only updates a local subgraph relevant to the current observation.
We define the retrieval ratio as the proportion of relationships updated at each iteration relative to the total number of relationships in the KG. As shown in Figure 8, this ratio (illustrated by gray bars) decreases steadily from 76% initially to 28% at iteration 20. This trend indicates that, as the KG grows, each update affects a progressively smaller fraction of the entire graph. This demonstrates that our spatial KG update mechanism effectively localizes modifications, ensuring computational efficiency and mitigating interference through context-aware incremental updates.
Figure 8: Average relationships related to update in spatial memory in each step.
Appendix C Dynamic Spatial Memory Update Algorithm
C.1 Detailed Algorithm of Spatial KG Update
0: New spatial knowledge graph $G_{\text{new}}=(V_{\text{new}},E_{\text{new}})$, main spatial knowledge graph $G=(V,E)$, queries $q\in Q$, entity & query embeddings $\mathcal{E}:V\cup Q\rightarrow\mathbb{R}^{d}$, maximum number of retrieved vertices $n$, maximum k hops $k$, vlm-base conflict resolver $\text{ResolveConflict}(\cdot)$
0: Updated consistent knowledge graph $G^{\prime}$
[…]
9: Remove isolated vertices from $G^{\prime}$
10:return $G^{\prime}$
Spatial Memory is a dynamically updated KG-based module designed to overcome agents’ limitations in spatial reasoning. Specifically, our Spatial Memory is formulated as a directed KG $G=(V,E)$, where $V$ denotes the set of all objects in the environment. Each object is a vertex in the KG. Each object’s name is encoded into a single semantic embedding vector via a pretrained embedding model [45]. The edge set $E$ captures spatial relationships between objects, each represented as a triple (e.g., $obj_{1},\text{relationship},obj_{2}]$). To update the Spatial KG, we need to continuously extract new relationships from current observations and update the old relationships in the Spatial KG. We use a VLM-based conflict resolver to address this problem. However, the more relationships provided to the conflict resolver, the more time it needs to update the KG. So we need to update relationships that are only related to the current situation. We design an algorithm that retrieves a sub-graph of KG that includes all vertices related to the current situation and both old and new relationships among them. We provide the sub-graph and new relationships to the conflict resolver. We need the conflict resolver to update the sub-graph based on information from the new relationships. The algorithm is shown in Algorithm 1.
To update $G$, we make use of the information provided by the step summarizer and query generator from the information preprocessor introduced in Sec. 2.1. First, a pretrained VLM-based Relation Retriever extracts the latest spatial relationships $G_{\text{new}}=(V_{\text{new}},E_{\text{new}})$ from the information provided by the step summarizer, which records high-level information in the current observation. Next, natural language queries (represented as $q\in Q$) (provided by the query generator) are used to retrieve relevant object vertices from $G$ via cosine similarity search. We select the top $n$ similar vertices compared with the query. These vertices are represent by $V_{\text{similar}}$.
Spatial KG maintains the relationships among different objects, so if we want to retrieve spatial information from spatial KG, we need to search for other objects that are related to the objects we observed in the current observation. In this way, we not only remember objects we can see, but also know the spatial information of the objects we cannot see. So we choose the k-hop algorithm to expand $V_{\text{similar}}$ using a K-hop neighborhood algorithm to capture contextually related objects. The K-hop algorithm is represented as $\text{K-hop}_{k}(V,G)$, which returns all vertices reachable within $\leq k$ hops from any vertices in $V$. The retrieved vertices are $V_{\text{expand}}$. We combine the vertices retrieved by cosine similarity and their K-hop neighbors to $V_{\text{retrieved}}$.
However, we need to resolve the conflict between new and old relationships. So $V_{\text{retrieved}}$ and the relationships among $V_{\text{retrieved}}$ is not enough. We need new vertices and relationships involved in the graph we provided to the VLM-based conflict resolver. To extract all vertices and relationships for the VLM-based conflict resolver and relationships, we not only need the relationships from KG (old information) and $G_{\text{new}}$, which represent new information. We need to connect old information and new information. To achieve this goal, we merge $G_{\text{new}}$ to $G$, which add new edges and vertices to $G$. We denote the merged KG as $G_{\text{union}}$. In $G_{\text{union}}$, we mix out-of-date and latest information. Then, we extract an induced sub-graph of $V_{\text{merged}}=V_{\text{retrieved}}\cup V_{\text{new}}$ from $G_{\text{union}}$. As both vertices from old graph $G$ and new graph is mixed in $V_{\text{merged}}$ and both edges from old and new graph is in $G_{\text{union}}$, the retrieved induced sub-graph $G_{\text{local}}$ contains all out-of-date and latest relationships among vertices that is related to current situation.
As $G_{\text{local}}$ contains all out-of-date and latest relationships and those relationships may have conflicts, a VLM-based conflict resolver (represent as $\text{ResolveConflict}(\cdot)$) is designed to resolve conflicts in $G_{\text{local}}$, and make sure that the relationship is the latest. The conflict resolver will take in $G_{\text{new}}$ and $G_{\text{local}}$, where $G_{\text{local}}$ is the graph waiting for update and $G_{\text{new}}$ provide update signal. The VLM-based conflict resolver will perform necessary updates such as adding vertices or inserting, deleting, or modifying relationships in $G_{\text{local}}$ based on $G_{\text{new}}$. The reconciled subgraph is then merged back into $G$, and any vertices that have lost all connections to other vertices during the update are pruned.
This design offers two key advantages: (1) Efficiency via localized updates: By restricting modifications to a context-relevant subgraph, we significantly reduce the number of relationships processed per update. Since VLMs struggle with reasoning over large sets of relationships, this constraint substantially improves both the efficiency and effectiveness of VLM-based KG updates. (2) Dynamic adaptability: The system continuously maintains up-to-date spatial knowledge, enabling agents to operate robustly in dynamic real-world environments.
Figure 9: Visualization of Spatial Memory’s dynamic update process.
C.2 Example of Dynamic Spatial Memory Update process
In RoboMemory’s Spatial Memory, the KG is dynamically constructed during environment exploration. As illustrated in Figure 9, we demonstrate the progressive expansion of the KG in Spatial Memory as the agent navigates through the environment. The figure indicates a continuous growth in the number of both vertices and edges of the KG as exploration progresses.
C.3 Proof of Dynamic Spatial Memory Update Algorithm
Theorem 1 (Upper Bound on K-hop Vertex Extraction in Directed Graphs).
Let $G=(V,E)$ be a finite directed graph with maximum out-degree $D\geq 1$, and let $\mathcal{S}\subseteq V$ be a set of $M$ source vertices. Define the K-hop neighborhood $\mathcal{N}_{K}(s)$ of a vertex $s\in\mathcal{S}$ as the set of vertices reachable from $s$ via directed paths of length at most $K$. Then the total number of distinct vertices in the union of all K-hop neighborhoods,
$\mathcal{N}_{K}(\mathcal{S})=\bigcup_{s\in\mathcal{S}}\mathcal{N}_{K}(s),$
Satisfies the following upper bound:
Proof.
For any vertex $s\in\mathcal{S}$, the number of distinct vertices reachable from $s$ within $i$ hops is at most $D^{i}$, assuming the worst-case scenario where each vertex encountered has the maximum out-degree $D$, and all neighbors are distinct and non-overlapping.
Thus, the size of the K-hop neighborhood of a single vertex satisfies:
Since there are $M$ such source vertices and assuming no overlaps between their K-hop neighborhoods (worst case), the union size satisfies:
$|\mathcal{N}_{K}(\mathcal{S})|\leq M\cdot|\mathcal{N}_{K}(s)|.$
Substituting the bound on $|\mathcal{N}_{K}(s)|$ gives the result. ∎
Theorem 2 (Upper Bound for K-hop Vertex Extraction in Normalized Directed Graphs).
Let $G=(V,E)$ be a finite directed graph with $|V|=n$ vertices. Assume the maximum out-degree is at most $D_{\max}=Dn$, and the maximum in-degree is at most $N_{\max}=Nn$, where $D,N\in(0,1]$ are constants. Let $\mathcal{S}\subseteq V$ be a set of $M$ source vertices. Define $\mathcal{N}_{K}(\mathcal{S})$ as the union of all vertices reachable from $\mathcal{S}$ via paths of length at most $K$, using only outgoing edges. Then the number of extracted vertices satisfies:
In particular, when $Dn\gg 1$, we have the approximation:
$|\mathcal{N}_{K}(\mathcal{S})|\lessapprox M\cdot(Dn)^{K}.$
Proof.
For each vertex $s\in\mathcal{S}$, the maximum number of reachable vertices within $i$-hops is at most $(Dn)^{i}$ under the assumption of maximum out-degree and no overlap.
Summing over hops from 0 to $K$, we get for each root:
$|\mathcal{N}_{K}(s)|\leq\sum_{i=0}^{K}(Dn)^{i}=\frac{(Dn)^{K+1}-1}{Dn-1}.$
Assuming no overlap among the $M$ source vertex expansions (worst case), we have:
$|\mathcal{N}_{K}(\mathcal{S})|\leq M\cdot\frac{(Dn)^{K+1}-1}{Dn-1}.$
Since the total number of vertices in the graph is $n$, this quantity is also trivially bounded above by $n$, yielding the result. ∎
Appendix D Additional Environment settings
D.1 EB-ALFRED and EB-Habitat
We adopt the same environment parameters as in EmbodiedBench. The maximum steps per task are set to 30, with image inputs of size 500 $\times$ 500. The temporal memory buffer length is set to 3. However, we modified the action formats of EB-ALFRED and EB-Habitat to simulate real-world scenarios better. Specifically, we define different action APIs (Python functions), where each action takes an object parameter indicating its target. We extract all possible objects from the environment as inputs to the Agent. The Agent must select appropriate actions and object parameters based on task requirements. Compared to the original interaction method in EmbodiedBench (which enumerates all possible actions, including both action names and target objects, and requires the Agent to choose), our approach offers greater flexibility. The detailed action APIs are presented in Table 4.
D.2 Real-world experiments